A stock has an expected return of 11 percent, its beta is 0.6, and the risk-free rate is 5.5 percent. What must the expected return on the market be?

15.40%

13.93%

14.67%

15.25%

9.17%
Answer #1

expected return=risk-free rate +Beta*(market rate- risk-free rate )

11=5.5+0.6*(Market rate-5.5)

(11-5.5)=0.6*(Market rate-5.5)

Market rate=(11-5.5)/0.6+5.5

=14.67%(Approx).